# Representation Learning of Lab Values via Masked AutoEncoder

### v2

@@ Abstract
Accurate imputation of missing laboratory values in electronic health records (EHRs) is critical to enable robust clinical predictions and reduce biases in AI systems in healthcare. Existing methods, such as variational autoencoders (VAEs) and decision tree-based approaches such as XGBoost, struggle to model the complex temporal and contextual dependencies in EHR data, mainly in underrepresented groups. In this work, we propose Lab-MAE, a novel transformer-based masked autoencoder framework that leverages self-supervised learning for the imputation of continuous sequential lab values. Lab-MAE introduces a structured encoding scheme that jointly models laboratory test values and their corresponding timestamps, enabling explicit capturing temporal dependencies. Empirical evaluation on the MIMIC-IV dataset demonstrates that Lab-MAE significantly outperforms the state-of-the-art baselines such as XGBoost across multiple metrics, including root mean square error (RMSE), R-squared (R2), and Wasserstein distance (WD). Notably, Lab-MAE achieves equitable performance across demographic groups of patients, advancing fairness in clinical predictions. We further investigate the role of follow-up laboratory values as potential shortcut features, revealing Lab-MAE’s robustness in scenarios where such data is unavailable. The findings suggest that our transformer-based architecture, adapted to the characteristics of the EHR data, offers a foundation model for more accurate and fair clinical imputation models. In addition, we measure and compare the carbon footprint of Lab-MAE with the baseline XGBoost model, highlighting its environmental requirements.
@@
The code to reproduce our experiments is available in github. The data set used is also publicly available [^1] This paper uses the MIMIC-IV dataset (Johnson et al., 2023), which is available in the PhysioNet repository (Johnson et al., 2020).
@@ 1 Introduction
Conventional imputation techniques, such as mean and standard deviation-based substitutions, are not well-suited to clinical tasks due to highly contextualized and individualized interpretation. These simplistic approaches fail to capture the intricate temporal and inter-variable dependencies present in high-dimensional physiological data Li et al. (2021). For example, the clinical importance of high Creatinine values in the hospitalized setting depends heavily on the patient’s own baseline values and the presence of any of the potential causes of acute kidney injury, including sepsis, hemorrhage, iatrogenic causes, urinary tract obstruction, and more. Some advanced methods, such as training multiple tabular models such as XGBoost Chen and Guestrin (2016) for individual lab values, have been used to capture these data relations in lab values Zhang et al. (2020); Chen and Guestrin (2016). However, even these models often struggle to fully leverage the available information, leading to suboptimal solutions that may overlook valuable contextual details Waljee et al. (2013); Luo et al. (2016).
@@
    We demonstrate that our model not only outperforms state-of-the-art tree-based methods, such as XGBoost, in the context of clinical data imputation but also mitigates biases by providing consistent performance across different patient demographics.
@@ 2.5.1 Baseline Models
To benchmark the performance of our proposed foundation Lab-MAE model for lab value imputation, we implemented a set of baseline models using XGBoost. XGBoost was selected for its well-known capabilities in handling tabular data tasks, including its ability to directly manage missing values without requiring extensive preprocessing. A total of 100 separate XGBoost models, value were trained, one for each lab value, to provide a robust comparison against our foundation Lab-MAE model.
@@
These trained XGBoost models served as our baseline for comparison with the Lab-MAE model. The key objective was to leverage XGBoost’s capability in handling missing data and its structured tree-based learning mechanism to predict individual lab values based on the contextual data.
2.5.2 Foundation Lab-MAE Model vs. XGB Evaluation Setup
Both the Lab-MAE and XGBoost models were evaluated using a cohort of 100,000 data points extracted from our independent test set. To ensure a fair comparison, the same data points were used across both models during evaluation. We computed three primary metrics for each lab value: RMSE, R2, and the WD. The R2 metric was used as the main reference due to its ability to assess the correlation strength and its capacity to avoid overfitting to mean or most common values. The WD was used to assess how well the model captures the overall distribution of lab values, including extreme values, unlike RMSE and R2, which emphasize overall accuracy and variance explained. WD complements RMSE and R2 by highlighting the quality of predictions in capturing not only central tendencies but also the full range of values, particularly the extremes. More information about the metrics is avaiable in appendix A.
@@ 2.5.3 Test Set Evaluation Process
Similarly, the XGBoost models were applied in a lab-specific manner, where each lab’s missing values were predicted using the corresponding XGBoost model trained for that specific lab. This approach allowed the XGBoost models to leverage their tree-based structure to effectively utilize the available data for predictions.
@@ 2.6.1 Lab-MAE Model Fairness Across Race Groups
To ensure that the Lab-MAE model’s predictions are equitable across different racial groups, we conducted an analysis of its performance for five race groups: White, Black, Hispanic, Asian, and Others. We compared the performance of the Lab-MAE model with the baseline XGBoost models for each race group using the same metrics.
@@ 2.6.2 Carbon Footprint Measurement
We assessed the carbon footprint of the Lab-MAE and baseline XGBoost models using the CodeCarbon library Courty et al. (2024). Emissions were estimated during the inference process for batch sizes of 1, 32, and 64 across multiple geographic locations, including Colombia, USA, France, Uganda, Philippines, and Australia. These locations were chosen to represent diverse geographic profiles in Asia, Africa, North America, South America, Europe, and Australia.
@@ 3 Results
3.1 Comparison of Lab-MAE vs. XGBoost for Data Imputation
In this section, we present the results of our comparison between the Lab-MAE and XGBoost models for the task of data imputation, evaluated using three metrics: WD, RMSE, and R2. These metrics were calculated on the entire set of 100 lab values and also on the top 20 most frequently occurring lab values to provide a more focused analysis.
@@ 3.1.1 Overall Analysis of All Lab Values
The comparison between the Lab-MAE and XGBoost models across all 100 lab values demonstrates that the Lab-MAE generally outperforms XGBoost. The average performance metrics calculated over the entire dataset indicate a lower WD and RMSE for the Lab-MAE model, and higher R2 values, suggesting that the Lab-MAE model has a better predictive capability and is less prone to error.
Table 1: Comparison of Lab-MAE and XGBoost across evaluation metrics for the 100 lab values. Each column indicates the amount of values the model is outperforming, where 100 means that a model outperforms in all the lab values for certain metric.
Metric Better Performer Better Performer
Lab-MAE (%) XGBoost (%)
RMSE 89 11
R2 89 11
Wasserstein Distance 79 21
As shown in Table 1, the comparison of Lab-MAE and XGBoost models across individual lab values provides a more nuanced evaluation of their performance. Instead of relying on mean values, which can obscure differences due to varying scales across lab tests, we analyze the number of lab values where each model performs better.
For WD, which evaluates the similarity between predicted and actual distributions (lower is better), Lab-MAE outperforms XGBoost in 79 out of 100 cases, demonstrating its superior ability to capture the underlying distributions, including extreme values.
For RMSE and R2, Lab-MAE achieves better performance in 89 cases, highlighting its robustness in minimizing overall error magnitude and better variance-capturing capabilities compared to XGBoost which outperforms Lab-MAE only in 11 cases.
Although XGBoost managed to surpass Lab-MAE in 11 cases for RMSE and R2, and 21 cases for WD, these instances were relatively isolated. This suggests that while XGBoost can occasionally outperform Lab-MAE in specific scenarios, the Lab-MAE model exhibits more stable and reliable performance across a broader range of laboratory values.
@@ 3.1.2 Focused Analysis on the Top 20 Lab Values
To provide a clearer understanding of the models’ performance on the most relevant lab values, we analyzed the top 20 most frequently occurring lab values. This focused analysis highlights the significant differences in prediction accuracy between Lab-MAE and XGBoost for these high-impact features.
@@ 4.1 Clinical and Technical Implications
Our LAB-MAE demonstrates a greater ability to learn real-world distributions, which may be multipolar or characterized by extreme values. This is highlighted by consistent R2 above XGB but reduced EMD/Wasserstein values overall and across subgroups. The model’s performance stability across demographic groups challenges the expected trade-off between accuracy and fairness in machine learning systems. This suggests that architectural innovations focused on capturing temporal and contextual relationships can simultaneously advance both objectives.
Another critical advantage of Lab-MAE lies in its design as a single foundation model capable of processing complex sequences of data, as opposed to the XGBoost approach that requires a separate model for each feature. This structural difference not only contributes to an improved energy efficiency, but also makes it a more scalable and practical solution for real-world applications.
@@ 4.2 Fairness and Equity Considerations
Previous approaches to laboratory value imputation have largely relied on traditional machine learning methods or simplified time-series models. Our model builds upon recent work by Bellamy et al. (2023) for modeling laboratory data, and Du et al. (2023) with the ReMasker framework for tabular data imputation, while addressing their limitations, particularly in handling temporal dependencies and maintaining performance across diverse patient populations. The significant improvement over XGBoost, especially in complex laboratory parameters, aligns with emerging evidence that properly architected deep learning models can overcome the traditional advantages of tree-based methods in tabular data van Breugel and van der Schaar (2024). However, the observed disparities in performance across racial groups reflect the social patterning of data generation Teotia et al. (2024), where factors such as systemic inequalities in healthcare access contribute to missingness patterns. Addressing this requires not only technical innovations in imputation models but also systemic efforts to improve data equity
@@ 4.3 Connection to Foundation Models
By consolidating multiple tasks into a single, cohesive model, Lab-MAE reduces the redundancy and overhead associated with training and deploying separate models for each feature, as required by XGBoost. This capability is particularly advantageous in healthcare settings like hospitals, where complex data environments demand robust yet streamlined solutions. Lab-MAE’s foundation model design highlights its potential to adapt effectively to diverse clinical scenarios, offering a powerful combination of scalability and environmental responsibility.
@@ 4.4 Limitations and Future Directions
Despite Lab-MAE’s promising results, several limitations merit attention. First, our evaluation was conducted on a single, albeit large, academic medical center dataset, potentially limiting generalizability. Second, while we demonstrated fairness across major demographic groups, future work should investigate intersectional fairness and rare subpopulations. Key directions for future research include: (1) extending Lab-MAE to incorporate structured medical knowledge, (2) investigating transfer learning capabilities across different healthcare settings, and (3) developing interpretability methods specifically designed for temporal clinical predictions.
@@ 5 Conclusion
[…]
Acknowledgments Yugang Jia is also affiliated with Verily Life Science LLC.

### v3

@@ Abstract
Accurate imputation of missing laboratory values in electronic health records (EHRs) is critical to enable robust clinical predictions and reduce biases in AI systems in healthcare. Existing methods, such as XGBoost, softimpute, GAIN, Expectation Maximization (EM), and MICE, struggle to model the complex temporal and contextual dependencies in EHR data, particularly in underrepresented groups. In this work, we propose Lab-MAE, a novel transformer-based masked autoencoder framework that leverages self-supervised learning for the imputation of continuous sequential lab values. Lab-MAE introduces a structured encoding scheme that jointly models laboratory test values and their corresponding timestamps, enabling explicit capturing temporal dependencies. Empirical evaluation on the MIMIC-IV dataset demonstrates that Lab-MAE significantly outperforms state-of-the-art baselines such as XGBoost, softimpute, GAIN, EM, and MICE across multiple metrics, including root mean square error (RMSE), R-squared (R2), and Wasserstein distance (WD). Notably, Lab-MAE achieves equitable performance across demographic groups of patients, advancing fairness in clinical predictions. We further investigate the role of follow-up laboratory values as potential shortcut features, revealing Lab-MAE’s robustness in scenarios where such data is unavailable. The findings suggest that our transformer-based architecture, adapted to the characteristics of EHR data, offers a foundation model for more accurate and fair clinical imputation. In addition, we measure and compare the carbon footprint of Lab-MAE with the a XGBoost model, highlighting its environmental requirements.
@@
The code to reproduce our experiments is available in anonymous github. The data set used is also publicly available [^1] This paper uses the MIMIC-IV dataset (Johnson et al., 2023), which is available in the PhysioNet repository (Johnson et al., 2020).
@@ 1 Introduction
Conventional imputation techniques, such as mean and standard deviation-based substitutions, are not well-suited to clinical tasks due to need of highly contextualized and individualized interpretation over complex data interactions. Other common imputation methods such as Softimpute Hastie et al. (2015); Platias and Petasis (2020), Expectation Maximization (EM) Li et al. (2021), and Multivariate Imputation by Chained-Equations (MICE) Li et al. (2021), fail to capture the intricate temporal and inter-variable dependencies present in high-dimensional physiological data Li et al. (2021). For example, the clinical importance of high Creatinine values in the hospitalized setting depends heavily on the patient’s own baseline values and the presence of any of the potential causes of acute kidney injury, including sepsis, hemorrhage, iatrogenic causes, urinary tract obstruction, and more. Some advanced methods to training multiple tabular models such as XGBoost Chen and Guestrin (2016) or even deep learning methods like GAN based methods such as GAIN Yoon et al. (2018); Li et al. (2021) for individual lab values, have been used to capture these data relations in lab values Zhang et al. (2020); Chen and Guestrin (2016). However, even these models often struggle to fully leverage the available information, leading to suboptimal solutions that may overlook valuable contextual details Waljee et al. (2013); Luo et al. (2016).
@@
    We demonstrate that our model not only outperforms state-of-the-art imputation methods, such as XGBoost, softimpute, GAIN, EM, and MICE, in the context of clinical data imputation but also mitigates biases by providing consistent performance across different patient demographics.
@@
Generalizable Insights about Machine Learning in the Context of Healthcare
Our work illustrates that a single, foundation-style Transformer model can simultaneously tackle multiple challenges—missing data, bias, and efficiency—within complex clinical datasets. In particular:
  • Unified Architecture for Tabular Data: Moving from separate, feature-specific models (e.g., one XGBoost per lab) to a single masked autoencoder not only streamlines deployment but also consistently improves imputation performance. This shift underscores the value of foundation-model thinking for healthcare’s heterogeneous, high-dimensional data.
  • Fairness Isn’t Necessarily a Trade-Off: Our findings suggest that high accuracy and equitable performance can be pursued together. By explicitly modeling temporal and contextual factors, we observed stable gains across diverse subgroups, highlighting that fairness can be integrated into design rather than added post hoc.
  • Efficiency and Carbon Footprint Matter: As ML models grow more complex, measuring and minimizing emissions becomes crucial—especially in time- and resource-constrained clinical settings. Our approach shows that a carefully architected, single-model strategy can be more environmentally sustainable than multiple, separate predictors.
@@ 2.5.1 Baseline Models
To benchmark the performance of our proposed Lab-MAE model for lab value imputation, we implemented a set of baseline models using XGBoost, softimpute, GAIN, EM, and MICE. For XGBoost implementation, a total of 100 separate XGBoost models, were trained, one for each lab value. For softimpute, GAIN, EM, and MICE, the models were implemented using hyperimpute Jarrett et al. (2022).
@@
2.5.2 Lab-MAE Imputation Evaluation Setup
Both the Lab-MAE and baseline models (XGBoost, softimpute, GAIN, EM, and MICE) were evaluated using a cohort of 100,000 data points extracted from our independent test set. To ensure a fair comparison, the same data points were used across both models during evaluation. We computed three primary metrics for each lab value: RMSE, R2, and the WD. The R2 metric was used as the main reference due to its ability to assess the correlation strength and its capacity to avoid overfitting to mean or most common values. The WD was used to assess how well the model captures the overall distribution of lab values, including extreme values, unlike RMSE and R2, which emphasize overall accuracy and variance explained. WD complements RMSE and R2 by highlighting the quality of predictions in capturing not only central tendencies but also the full range of values, particularly the extremes. More information about the metrics is avaiable in appendix A.
@@ 2.5.3 Test Set Evaluation Process
Similarly, the baseline methods were trained in a lab-specific manner, where each lab’s missing values were predicted using each imputation method.
@@ 2.6.1 Lab-MAE Model Fairness Across Race Groups
To ensure that the Lab-MAE model’s predictions are equitable across different racial groups, we conducted an analysis of its performance for five race groups: White, Black, Hispanic, Asian, and Others. We compared the performance of the Lab-MAE model with the best-performed baseline model (XGBoost) for each race group using the same metrics.
@@ 2.6.2 Carbon Footprint Measurement
We assessed the carbon footprint of the Lab-MAE and the best-performed baseline (XGBoost), using the CodeCarbon library Courty et al. (2024). Emissions were estimated during the inference process for batch sizes of 1, 32, and 64 across multiple geographic locations, including Colombia, USA, France, Uganda, Philippines, and Australia. These locations were chosen to represent diverse geographic profiles in Asia, Africa, North America, South America, Europe, and Australia.
@@ 3 Results
3.1 Lab-MAE imputation performance
In this section, we present the results of our comparison between the Lab-MAE and the baseline models (XGBoost, softimpute, GAIN, EM, and MICE) for the task of data imputation. We evaluated using three metrics: WD, RMSE, and R2. These metrics were calculated on the entire set of 100 lab values. A further detailed comparison with XGBoost on the top 20 most frequently occurring lab values was performed to provide more insights about model performance.
@@ 3.1.1 Overall Analysis of All Lab Values
To mitigate biases arising from differences in the scale of lab measurements and the inherent imbalance in lab value frequencies, we aggregated the evaluation metrics across 100 lab tests by counting the number of tests in which each baseline method outperforms Lab-MAE. In this analysis, for RMSE and WD, lower values indicate superior performance, whereas for R2 higher values are preferable. This aggregation strategy provides a more balanced assessment of performance across heterogeneous lab tests.
Table 1: Comparison of Lab-MAE with baseline imputation methods across evaluation metrics. The figures represent the number of lab tests (out of 100) where the baseline method outperforms Lab-MAE. For RMSE and WD, lower values indicate better performance, whereas for R2, higher values are preferred.
Model vs Ours RMSE R2 Wasserstein Distance
XGBoost 11% 11% 21%
softimpute 14% 0% 6%
GAIN 3% 5% 11%
EM 4% 12% 12%
MICE 11% 18% 16%
Table 1 summarizes the comparison between Lab-MAE and the baseline imputation methods—XGBoost, softimpute, GAIN, EM, and MICE. For example, when compared to XGBoost, Lab-MAE achieves lower RMSE and higher R2 in 89 out of 100 lab tests, and a lower WD in 79 tests. In contrast, the other baselines outperform Lab-MAE in fewer tests: softimpute surpasses Lab-MAE in 14 tests for RMSE, 6 for WD, and in none for R2; GAIN outperforms in 3 (RMSE), 11 (WD), and 5 (R2) tests; EM does so in 4 (RMSE), 12 (WD), and 12 (R2) tests; and MICE in 11 (RMSE), 16 (WD), and 18 (R2) tests.
@@ 3.1.2 Focused Analysis on the Top 20 Lab Values
To provide a clearer understanding of the models’ performance on the most relevant lab values, we analyzed the top 20 most frequently occurring lab values. This focused analysis highlights the differences in lab value imputaion between Lab-MAE and the best-performed model-XGBoost.
@@ 4.1 Clinical and Technical Implications
Our LAB-MAE demonstrates a greater ability to learn real-world distributions, which may be characterized by extreme values. This is highlighted by consistent R2 above baselines but reduced EMD/Wasserstein values overall and across subgroups.
The model’s performance stability across demographic groups challenges the expected trade-off between accuracy and fairness in machine learning systems. This suggests that architectural innovations focused on capturing temporal and contextual relationships can simultaneously advance both objectives.
@@ 4.2 Fairness and Equity Considerations
Previous approaches to laboratory value imputation have largely relied on traditional machine learning methods or simplified time-series models. Our model builds upon recent work by Bellamy et al. (2023) for modeling laboratory data, and Du et al. (2023) with the ReMasker framework for tabular data imputation, while addressing their limitations, particularly in handling temporal dependencies and maintaining performance across diverse patient populations. The significant improvement over the baseline models, especially in complex laboratory parameters, aligns with emerging evidence that properly architected deep learning models can overcome the traditional advantages of tree-based methods in tabular data van Breugel and van der Schaar (2024). However, the observed disparities in performance across racial groups reflect the social patterning of data generation Teotia et al. (2024), where factors such as systemic inequalities in healthcare access contribute to missingness patterns. Addressing this requires not only technical innovations in imputation models but also systemic efforts to improve data equity
@@ 4.3 Connection to Foundation Models
By consolidating multiple tasks into a single, cohesive model, Lab-MAE reduces the redundancy and overhead associated with training and deploying separate models for each feature, as required by XGBoost or MICS. This capability is particularly advantageous in healthcare settings like hospitals, where complex data environments demand robust yet streamlined solutions. Lab-MAE’s foundation model design highlights its potential to adapt effectively to diverse clinical scenarios, offering a powerful combination of scalability and environmental responsibility.
@@ 4.4 Limitations and Future Directions
Despite Lab-MAE’s promising results, several limitations merit attention. First, our evaluation was conducted on a single, academic medical center dataset, potentially limiting generalizability. Second, while we demonstrated fairness across major demographic groups, future work should investigate intersectional fairness and rare subpopulations. Key directions for future research include: (1) extending Lab-MAE to incorporate structured medical knowledge, (2) investigating transfer learning capabilities across different healthcare settings, and (3) developing interpretability methods specifically designed for temporal clinical predictions.